# AV-RIR: Audio-Visual Room Impulse Response Estimation

### v1

@@ Abstract
Accurate estimation of Room Impulse Response (RIR), which captures an environment’s acoustic properties, is important for speech processing and AR/VR applications. We propose AV-RIR, a novel multi-modal multi-task learning approach to accurately estimate the RIR from a given reverberant speech signal and the visual cues of its corresponding environment. AV-RIR builds on a novel neural codec-based architecture that effectively captures environment geometry and materials properties and solves speech dereverberation as an auxiliary task by using multi-task learning. We also propose Geo-Mat features that augment material information into visual cues and CRIP that improves late reverberation components in the estimated RIR via image-to-RIR retrieval by 86%. Empirical results show that AV-RIR quantitatively outperforms previous audio-only and visual-only approaches by achieving 36% - 63% improvement across various acoustic metrics in RIR estimation. Additionally, it also achieves higher preference scores in human evaluation. As an auxiliary benefit, dereverbed speech from AV-RIR shows competitive performance with the state-of-the-art in various spoken language processing tasks and outperforms reverberation time error score in the real-world AVSpeech dataset. Qualitative examples of both synthesized reverberant speech and enhanced speech are available online [^1] https://www.youtube.com/watch?v=tTsKhviukAE.
@@ 1 Introduction
Main Contributions. We propose AV-RIR, a novel multi-modal multi-task learning approach for RIR estimation. We present a novel neural codec-based multi-modal architecture that takes as input a reverberant speech uttered in a source environment, the panoramic image of the environment, and our novel Geo-Mat feature that incorporates information about room geometry and the materials of surfaces and objects. Our model is equipped with carefully designed encoders, decoders, and a Residual Vector Quantizer that learns rich task-specific features while discarding the noise in training data [94]. Additionally, the model incorporates a dual-branch structure for multi-task learning, where we solve an auxiliary speech dereverberation task alongside the primary RIR estimation task. This approach effectively redefines the ultimate learning objective as decomposing reverberated speech into its constituent anechoic speech and RIR components. Furthermore, we propose Contrastive RIR-Image Pre-training (CRIP) to improve late reverberation of the estimated RIR during inference time using image-to-RIR retrieval. To summarize, our main contributions are as follows:
@@
    AV-RIR employs a neural codec-based multi-modal architecture that inputs audio and visual cues and a novel Geo-Mat feature that captures room geometry and material information. We also propose CRIP to improve late reverberation of the estimated RIR using retrieval and observe that CRIP improves late reverberation by 86%.
@@
    AV-RIR solves an auxiliary speech dereverberation task for learning RIR estimation. Through this, AV-RIR essentially learns to separate anechoic speech and RIR.
@@ 2 Background and Related Work
where RIR represents the intensity and time of arrival of direct sound, early reflections, and late reverberation. The RIR can either be measured in a controlled environment [79, 6, 25] or simulated using physics-based simulators [95, 88]. Measuring RIR requires sophisticated hardware and human labor. At the same time, acoustic simulators require a 3D mesh representation of the environment [87] and complete knowledge of room materials [78]. Thus, these simulators cannot capture all the acoustic effects of an RIR at an interactive rate [87]. Prior works also show that accurately generating RIRs without these cues is infeasible (e.g., only from RGB images [81, 55]). Alternatively, estimating RIR from reverberant speech has seen better success [83]. Reverberant speech can be easily recorded using household devices (e.g., mobile phone, Amazon Echo, etc.). Prior work explores audio-only techniques [71, 83, 45], and to the best of our knowledge, AV-RIR is the first audio-visual method to solve this task.
@@ 3.1 Overview: AV-RIR
Fig. 2 gives an overview of our approach. Given a reverberant speech $\mathcal{S}_{R}$, the task of AV-RIR is to learn to estimate the RIR from $\mathcal{S}_{R}$. To achieve this, we propose a novel multi-modal neural architecture and solve two parallel tasks for learning accurate RIR estimation. As input, together with the $\mathcal{S}_{R}$, AV-RIR also receives the RGB panoramic image $\mathcal{I}_{P}$ of the source environment and our proposed Geo-Mat feature map $\mathcal{I}_{G}$. The construction of $\mathcal{I}_{G}$ is illustrated in Fig. 3. The $\mathcal{S}_{R}$ is first passed through a reverberant speech encoder, after which AV-RIR breaks into two branches that solve two different tasks. The bottom branch, which also receives ResNet-18 encoded features of $\mathcal{I}_{G}$, solves our primary RIR task estimation. The other branch, which receives the encoded features of $\mathcal{I}_{P}$, solves an auxiliary speech dereverberation task to predict enhanced speech $\mathcal{S}_{C}$. After the multi-modal feature fusion step, both branches employ a Residual Vector Quantizer module. During inference, to synthesize any target speech as if spoken in the source environment, we convolve the estimated RIR with the target clean speech. Additionally, we propose CRIP (Fig. 4) to retrieve an RIR from a datastore $\mathcal{DS}$, conditioned on the $\mathcal{I}_{P}$ of the source environment, to improve late components. Next, we will describe each module in detail.
@@ 3.2 AV-RIR Architecture
Speech Dereverberation Encoder ($\mathcal{E}_{{S}}$). Our $\mathcal{E}_{{S}}$ follows the streaming SEANet encoder [47] without the skip connections proposed in SoundStream [96, 92]. We modified the architecture to have five convolutional blocks with 2, 2, 3, 4, and 5 stride and 2, 4, 8, 16, and 32 output channels, respectively. Following prior audio codec [93], our architecture effectively downsamples the speech signals by 240.
Room Impulse Response Encoder ($\mathcal{E}_{{IR}}$). Our $\mathcal{E}_{{IR}}$ is adapted and modified from the S2IR-GAN encoder [71]. Our three-layer $\mathcal{E}_{{IR}}$ has 256, 512, and 1024 output channels, 14401, 41, and 41 kernel lengths, and 225, 2, and 2 strides, respectively. The large kernel length in the first layer encodes the RIR features efficiently. We significantly reduce the input dimension by a factor of 900. We process reverberant speech $\mathcal{S}_{R}$ segments of $\mathbb{R}^{1\times 14400}$ samples. Therefore, every reverberant speech sample is encoded into $\mathbb{R}^{1024\times 16}$ RIR temporal features.
Vision Encoders ($\mathcal{E}_{{P}}$, $\mathcal{E}_{{GM}}$). Prior work on audio-visual speech dereverberation and localization has shown that ResNet-18 [30] is capable of extracting strong cues from image $\mathcal{I}_{P}$ and depth maps [16, 55, 15]. Therefore, we employ two separate pre-trained ResNet-18-based feature encoders $\mathcal{E}_{{P}}$ and $\mathcal{E}_{{GM}}$ to encode the $\mathcal{I}_{P}$ and the Geo-Mat feature $\mathcal{I}_{G}$, respectively, and reshape the features to $\mathbb{R}^{1024\times 4}$.
@@
Decoders ($\mathcal{D}_{IR}$, $\mathcal{D}_{S}$). As the two branches AV-RIR are responsible for decoding separate outputs from the compressed codes (i.e., enhanced speech and RIR), we use two different decoders for this task. We use a HiFi-GAN vocoder for the speech dereverberation branch to decode enhanced speech from the compressed code. HiFi-GAN [41] has shown impressive performance in generating high-fidelity speech, especially with audio codecs [92]. For the RIR estimation branch, we use a modified SoundStream decoder [96]. We modify the decoder to have 6 transposed convolutional (Conv) blocks with output channels of (256, 128, 64, 32, 32, 16) and strides of (5, 5, 2, 2, 1, 1). The output from the last transposed Conv block is passed to a final 1D Conv layer with kernel size 1 and stride 1 to project the code to the waveform domain.
@@ 3.3 Geo-Mat Features
Figure 3: The computation pipeline of Geo-Mat feature map. The first two channels of the Geo-Mat feature ($\mathcal{I}_{G}$) comprise the absorption coefficients ($\mathcal{AC}$) of each acoustic material. The third channel comprises the depth map. We illustrate objects in the environment having similar $\mathcal{AC}$ with similar colors: chairs and furniture with similar materials are represented in light blue, painting, and wall pictures with similar materials are represented in yellow, and the rest in grey. Identifying the objects and their $\mathcal{AC}$ involves a pipeline with several out-of-the-box pre-trained models and an acoustic material database, described in Section 3.3.
@@ 3.5 Contrastive RIR-Image Pretraining (CRIP)
We propose CRIP, a model built on the fundamentals of CLIP [64], that learns a joint embedding space between the panoramic image ($\mathcal{I}_{P}$) and their corresponding RIRs. Similar to CLIP, CRIP employs two encoders, a pre-trained HorizonNet encoder [85] $\mathcal{E}_{H}$ which serves as our $\mathcal{I}_{P}$ encoder, and the discriminator network proposed in FAST-RIR [70], which serves as our RIR encoder $\mathcal{E}_{F}$. Formally, $\mathcal{E}_{H}$ takes as input $\mathcal{I}_{P}$ and outputs an embedding $\mathcal{I}_{EM}\in\mathbb{R}^{N\times 1024}$, and $\mathcal{E}_{F}$ takes as input $\mathcal{RIR}\in\mathbb{R}^{N\times 3\times 1024}$ and output an embedding $\mathcal{R}_{EM}\in\mathbb{R}^{N\times 1024}$. Finally, we measure similarity by calculating the dot product between the two as follows:
$C_{r-2-i}=\tau*\left(\mathcal{RIR}\cdot\mathcal{I}_{EM}^{\top}\right)$ (10)
$C_{i-2-r}=\tau*\left(\mathcal{I}_{EM}\cdot\mathcal{RIR}^{\top}\right)$
where $\tau$ is the temperature. This is followed by calculating the RIR-to-Image loss $\ell_{r-2-i}$ and the Image-to-RIR loss $\ell_{r-2-i}$ as follows:
$\ell_{r-2-i}=\frac{1}{N}\sum_{i=0}^{N}\log\operatorname{diag}(\operatorname{softmax}(C_{r-2-i}))$ (11)
$\ell_{r-2-i}=\frac{1}{N}\sum_{i=0}^{N}\log\operatorname{diag}(\operatorname{softmax}(C_{i-2-r}))$
@@ 4.1 RIR Estimation
Baselines. We compare the performance of AV-RIR for RIR estimation with six other baselines. 1. Image2Reverb [81]: The Image2Reverb predicts the RIR from the camera-view image. 2. Visual Acoustic Matching (VAM) [13]: VAM inputs the source audio and the target environment image and outputs resynthesized audio matching the target environment. VAM does not explicitly estimate the RIR of the target environment. Therefore, we compare VAM only on our perceptual evaluation. 3. FAST-RIR++  [70]: FAST-RIR inputs the room geometry, source and listener positions, and $T_{60}$ to generate the RIR. We modified the architecture to the input panoramic image ($\mathcal{I}_{P}$) of the environment and estimate the RIR (FAST-RIR++). 4. CRIP-only (ours): We propose CRIP to retrieve the closest RIR from the large synthetic dataset $\mathcal{DS}$ using a panoramic image $\mathcal{I}_{P}$ as input. We use it as a baseline to evaluate how much improvement our audio-visual network contributes. 5. Filtered Noise Shaping Network (FiNS) [83]: FiNS is the audio-only time domain RIR estimator from reverberant speech. 6. S2IR-GAN  [71]: S2IR-GAN is an audio-only GAN-based reverberant speech-to-IR estimator.
@@
Speech ${}^{\dagger}$ Speaker ${}^{\dagger}$ RTE* ↓
@@ 4.2 Speech Dereverberation
Evaluation Metrics. We evaluate speech dereverberation using our AV-RIR on two downstream tasks: automatic speech recognition (ASR) and speaker verification (SV). We use standard metrics such as Word Error Rate (WER) to measure ASR performance and Equal Error Rate (EER) to measure the SV. Following prior works [16, 18], we used pre-trained models from SpeechBrain [72] to evaluate the ASR and SV performance. The recorded AVSpeech dataset does not have parallel ground-truth clean speech to perform ASR and SV tasks. Therefore, we used the Reverberation Time Error (RTE) metric proposed in Chowdhury et al. [18] to evaluate the estimated clean speech from our AV-RIR.
Baselines. We compared the speech dereverberation using our approach with the following prior audio-only and audio-visual speech enhancement networks. 1. Audio-Only: WPE [59] is a statistical-model-based speech dereverberation network that cancels late reverberation without the knowledge of RIR. MetricGan+ [27], DEMUCS [22], HiFi-GAN [84], VoiceFixer [51], SkipConvGAN [43] and Kotha et al. [42] are learning-based speech dereverberation networks. 2. Audio-Visual : VIDA [16] is the first audio-visual speech dereverberation network that takes $\mathcal{I}_{P}$ as visual cues. Recently, geometry-aware AdVerb [18] has been shown to achieve SOTA results in downstream speech tasks.
@@
Ablation Study. We also perform a comprehensive ablation study to show the benefit of different components of our AV-RIR. (1) Multi-task learning. To prove the effectiveness of the multi-task learning approach, we train the branches separately (AV-RIR w/o Multi-Task). Table 1 and Table 2 show that our multi-task learning approach benefits both tasks mutually. While the RIR estimation performance improves by 31% - 48%, speech dereverberation performance improves by 13% - 21%. (2) CRIP. Table 1 shows that adding CRIP during inference for late reverberation improves the late component MSE (LMSE) by 86%. (3) Geo-Mat feature. Table 1 shows that the Geo-Mat feature improves RIR estimation accuracy by 11% - 28%. (4) Visual cues. To prove the effectiveness of visual cues, we discard them while training and directly pass RIR and speech encoder outputs to RVQ. Table 1 and Table 2 show that AV-RIR outperforms our audio-only AV-RIR variation in RIR estimation tasks by 41% - 55% and speech dereverbation task by around 24%.

### v2

@@ Abstract
Accurate estimation of Room Impulse Response (RIR), which captures an environment’s acoustic properties, is important for speech processing and AR/VR applications. We propose AV-RIR, a novel multi-modal multi-task learning approach to accurately estimate the RIR from a given reverberant speech signal and the visual cues of its corresponding environment. AV-RIR builds on a novel neural codec-based architecture that effectively captures environment geometry and materials properties and solves speech dereverberation as an auxiliary task by using multi-task learning. We also propose Geo-Mat features that augment material information into visual cues and CRIP that improves late reverberation components in the estimated RIR via image-to-RIR retrieval by 86%. Empirical results show that AV-RIR quantitatively outperforms previous audio-only and visual-only approaches by achieving 36% - 63% improvement across various acoustic metrics in RIR estimation. Additionally, it also achieves higher preference scores in human evaluation. As an auxiliary benefit, dereverbed speech from AV-RIR shows competitive performance with the state-of-the-art in various spoken language processing tasks and outperforms reverberation time error score in the real-world AVSpeech dataset. Qualitative examples of both synthesized reverberant speech and enhanced speech are available online [^1] https://anton-jeran.github.io/AVRIR/.
@@ 1 Introduction
Main Contributions. We propose AV-RIR, a novel multi-modal multi-task learning approach for RIR estimation. AV-RIR employs a novel neural codec-based multi-modal architecture that takes as input a reverberant speech uttered in a source environment, the panoramic image of the environment, and a novel Geo-Mat feature that incorporates information about room geometry and the materials of surfaces and objects. The multi-modal architecture consists of carefully designed encoders, decoders, and a Residual Vector Quantizer that learns rich task-specific (RIR estimation and speech deverberation) features while discarding the noise in training data [94]. Additionally, AV-RIR incorporates a dual-branch structure for multi-task learning, where we solve an auxiliary speech dereverberation task alongside the primary RIR estimation task. This approach effectively redefines the ultimate learning objective as decomposing reverberated speech into its constituent anechoic speech and RIR components. Furthermore, we propose Contrastive RIR-Image Pre-training (CRIP) to improve late reverberation of the estimated RIR during inference time using image-to-RIR retrieval. To summarize, our main contributions are as follows:
@@
    AV-RIR employs a neural codec-based multi-modal architecture that takes as input audio, visual cues and a novel Geo-Mat feature. We also propose CRIP to improve late reverberation effects using retrieval.
@@
    During training, AV-RIR solves an auxiliary speech dereverberation task for learning RIR estimation. Through this, AV-RIR essentially learns to separate anechoic speech and RIR.
@@ 2 Background and Related Work
where RIR represents the intensity and time of arrival of direct sound, early reflections, and late reverberation. The RIR can either be measured in a controlled environment [79, 6, 25, chen2024RAF] or simulated using physics-based simulators [95, 88]. Measuring RIR requires sophisticated hardware and human labor. At the same time, acoustic simulators require a 3D mesh representation of the environment [87] and complete knowledge of room materials [78]. Thus, these simulators cannot capture all the acoustic effects of an RIR at an interactive rate [87]. Prior works also show that accurately generating RIRs without these cues is infeasible (e.g., only from RGB images [81, 55]). Alternatively, estimating RIR from reverberant speech has seen better success [83]. Reverberant speech can be easily recorded using household devices (e.g., mobile phone, Amazon Echo, etc.). Prior work explores audio-only techniques [71, 83, 45], and to the best of our knowledge, AV-RIR is the first audio-visual method to solve this task.
@@ 3.1 Overview: AV-RIR
Fig. 2 gives an overview of our approach. Given a reverberant speech $\mathcal{S}_{R}$, the task of AV-RIR is to learn to estimate the RIR from $\mathcal{S}_{R}$. To achieve this, we propose a novel multi-modal neural architecture and solve two parallel tasks for learning accurate RIR estimation. As input, together with the $\mathcal{S}_{R}$, AV-RIR also receives the RGB panoramic image $\mathcal{I}_{P}$ of the source environment and our proposed Geo-Mat feature map $\mathcal{I}_{G}$. The construction of $\mathcal{I}_{G}$ is illustrated in Fig. 3. The $\mathcal{S}_{R}$ is first passed through a reverberant speech encoder, after which AV-RIR breaks into two branches that solve two different tasks. The bottom branch, which also receives ResNet-18 encoded features of $\mathcal{I}_{G}$, solves our primary RIR estimation task. The other branch, which receives the encoded features of $\mathcal{I}_{P}$, solves an auxiliary speech dereverberation task to predict enhanced speech $\mathcal{S}_{C}$. After the multi-modal feature fusion step, both branches employ a Residual Vector Quantizer module. During inference, to synthesize any target speech as if spoken in the source environment, we convolve the estimated RIR with the target clean speech. Additionally, we propose CRIP (Fig. 4) to retrieve an RIR from a datastore $\mathcal{DS}$, conditioned on the $\mathcal{I}_{P}$ of the source environment, to improve late components. Next, we will describe each module in detail.
@@ 3.2 AV-RIR Architecture
Room Impulse Response Encoder ($\mathcal{E}_{{IR}}$). $\mathcal{E}_{{IR}}$ is adapted and modified from the S2IR-GAN encoder [71]. Our three-layer $\mathcal{E}_{{IR}}$ has 256, 512, and 1024 output channels, 14401, 41, and 41 kernel lengths, and 225, 2, and 2 strides, respectively. The large kernel length in the first layer encodes the RIR features efficiently. We significantly reduce the input dimension by a factor of 900. We process reverberant speech $\mathcal{S}_{R}$ segments of $\mathbb{R}^{1\times 14400}$ samples. Therefore, every reverberant speech sample is encoded into $\mathbb{R}^{1024\times 16}$ RIR temporal features.
Vision Encoders ($\mathcal{E}_{{P}}$, $\mathcal{E}_{{GM}}$). Prior work on audio-visual speech dereverberation and localization has shown that ResNet-18 [30] is capable of extracting strong cues from image $\mathcal{I}_{P}$ and depth maps [16, 55, 15]. Therefore, we employ two separate ResNet-18-based feature encoders $\mathcal{E}_{{P}}$ and $\mathcal{E}_{{GM}}$ to encode the $\mathcal{I}_{P}$ and the Geo-Mat feature $\mathcal{I}_{G}$, respectively, and reshape the features to $\mathbb{R}^{1024\times 4}$.
@@
Decoders ($\mathcal{D}_{IR}$, $\mathcal{D}_{S}$). As the two branches of AV-RIR are responsible for decoding separate outputs from the compressed codes (i.e., enhanced speech and RIR), we use two different decoders for this task. We use a HiFi-GAN vocoder for the speech dereverberation branch to decode enhanced speech from the compressed code. HiFi-GAN [41] has shown impressive performance in generating high-fidelity speech, especially with audio codecs [92]. For the RIR estimation branch, we use a modified SoundStream decoder [96]. We modify the decoder to have 6 transposed convolutional (Conv) blocks with output channels of (256, 128, 64, 32, 32, 16) and strides of (5, 5, 2, 2, 1, 1). The output from the last transposed Conv block is passed to a final 1D Conv layer with kernel size 1 and stride 1 to project the code to the waveform domain.
@@ 3.3 Geo-Mat Features
Figure 3: The computation pipeline of Geo-Mat feature map. The first two channels of the Geo-Mat feature ($\mathcal{I}_{G}$) comprise the absorption coefficients ($\mathcal{AC}$) of each acoustic material. The third channel comprises the depth map. We illustrate objects in the environment having similar $\mathcal{AC}$ with similar colors: chairs and furniture with similar materials are represented in light blue, painting, and wall pictures with similar materials are represented in yellow, and the rest in grey. More details on the method to obtain $\mathcal{AC}$ is described in Section 3.3.
@@ 3.5 Contrastive RIR-Image Pretraining (CRIP)
We propose CRIP, a model built on the fundamentals of CLIP [64], that learns a joint embedding space between the panoramic image ($\mathcal{I}_{P}$) and their corresponding RIRs. Similar to CLIP, CRIP employs two encoders, a pre-trained HorizonNet encoder [85] $\mathcal{E}_{H}$ which serves as our $\mathcal{I}_{P}$ encoder, and the discriminator network proposed in FAST-RIR [70], which serves as our RIR encoder $\mathcal{E}_{F}$. Formally, $\mathcal{E}_{H}$ takes as input $\mathcal{I}_{P}$ and outputs an embedding $\mathcal{I}_{EM}\in\mathbb{R}^{N\times 1024}$, and $\mathcal{E}_{F}$ takes as input $\mathcal{RIR}\in\mathbb{R}^{N\times 1\times 4096}$ and output an embedding $\mathcal{R}_{EM}\in\mathbb{R}^{N\times 1024}$. Finally, we measure similarity by calculating the dot product between the two as follows:
[…]
@@ 4.1 RIR Estimation
Baselines. We compare the performance of AV-RIR with six other baselines. (1) Image2Reverb [81]: Predicts RIR from the camera-view image. (2) Visual Acoustic Matching (VAM) [13]: Takes as input the source audio and the target environment image and outputs resynthesized audio matching the target environment. VAM does not explicitly estimate the RIR of the target environment. Therefore, we compare VAM only on our perceptual evaluation. (3) FAST-RIR++  [70]: Takes as input the room geometry, source and listener positions, and $T_{60}$ to generate the RIR. We modified the architecture to the input panoramic image ($\mathcal{I}_{P}$) of the environment and estimate the RIR (FAST-RIR++). (4) CRIP-only (ours): CRIP retrieves the closest RIR from the large synthetic dataset $\mathcal{DS}$ using a panoramic image $\mathcal{I}_{P}$ as input. We use it as a baseline to evaluate how much improvement our audio-visual network contributes. (5) Filtered Noise Shaping Network (FiNS) [83]: An audio-only time domain RIR estimator from reverberant speech. (6) S2IR-GAN  [71]: An audio-only GAN-based reverberant speech-to-IR estimator.
@@
Speech Recognition Speaker Verification RTE* ↓
@@ 4.2 Speech Dereverberation
Evaluation Metrics. We evaluate speech dereverberation using our AV-RIR on two downstream tasks: Automatic Speech Recognition (ASR) and Speaker Verification (SV). We use standard metrics such as Word Error Rate (WER) to measure ASR performance and Equal Error Rate (EER) to measure the SV. Following prior works [16, 18], we used pre-trained models from SpeechBrain [72] to evaluate the ASR and SV performance. AVSpeech dataset does not have parallel clean speech to perform ASR and SV tasks and we use the Reverberation Time Error (RTE) metric [18] to evaluate the estimated clean speech from our AV-RIR.
Baselines. We compared the speech dereverberation using our approach with the following prior audio-only and audio-visual speech enhancement networks. (1) Audio-Only: WPE [59] is a statistical-model-based speech dereverberation network that cancels late reverberation without the knowledge of RIR. MetricGan+ [27], DEMUCS [22], HiFi-GAN [84], VoiceFixer [51], SkipConvGAN [43] and Kotha et al. [42] are learning-based speech dereverberation networks. (2) Audio-Visual : VIDA [16] is the first audio-visual speech dereverberation network that takes $\mathcal{I}_{P}$ as visual cues. Recently, geometry-aware AdVerb [18] has been shown to achieve SOTA results in downstream speech tasks.
@@
Ablation Study. We perform a comprehensive ablation study to show the benefit of different components of our AV-RIR. (1) Multi-task learning. To prove the effectiveness of the multi-task learning approach, we train the branches separately (AV-RIR w/o Multi-Task). Table 1 and Table 2 show that our multi-task learning approach benefits both tasks mutually. While the RIR estimation performance improves by 31% - 48%, speech dereverberation performance improves by 13% - 21%. (2) CRIP. Table 1 shows that adding CRIP during inference for late reverberation improves the late component MSE (LMSE) by 86%. (3) Geo-Mat feature. Table 1 shows that the Geo-Mat feature improves RIR estimation accuracy by 11% - 28%. (4) Visual cues. To prove the effectiveness of visual cues, we discard them while training and directly pass RIR and speech encoder outputs to RVQ. Additionally, we also discard CRIP and directly estimate the full RIR. Table 1 and Table 2 show that AV-RIR outperforms our audio-only AV-RIR variation in RIR estimation tasks by 41% - 55% and speech dereverbation task by around 24%.